# KEN KANGDA WREN1

### v4

@@ B PROOF OF PROPOSITION 1
This has solutions of the form $\{A^{+}_{t}\}=\{\Pi^{+}_{t}\}$ only, where $\{\Pi^{+}_{t}\}$ is any model-risk inference based on any measure $\Pi^{B}_{0}W_{B}^{Z}|_{T}$, $\Pi^{B}_{0}\sim\pi^{B}_{0}$. The same can be obtained from the LHS of (B.5).
@@
Adding $B$-conditionally independent data $\{D_{t}\}$ under law $W_{B}^{D}|_{T}$, the argument above applies unchanged to the dynamics of $\{dS^{(0|A)}_{t}\}$, (B.1-B.2), under reference belief ${\pi}^{B}_{0}\mathbf{W}_{B}^{T}$ and some RNE measure $\hat{\pi}^{B}_{0}\hat{\mathbf{W}}_{B}^{T}$ such that (B.4) holds and $d(\hat{\mathbf{W}}_{B}^{T})_{t}=\hat{r}_{B,\text{ }t}dt+d(\mathbf{W}_{-}^{T})_{t}$, $\hat{{r}}_{B,\text{ }t}dt=-d(\hat{{RP}}^{B}_{t})$. To conclude, the RNE measure $\{S^{(0|A)}_{t}\}$ must have the canonical form: $\Pi^{B}_{0}\mathbf{W}_{B}^{T}$, $\Pi^{B}_{0}\sim\pi^{B}_{0}$.
Part II. Viability of canonical normal pricing $\{S^{(RP|{\Pi})}_{t}\}=\{{S}^{(0|{\Pi})}_{t}\}-\{\Sigma_{B}{RP}^{B}(t)\Pi^{B}_{t}\}$.
@@
Under the martingale measure $\Pi^{B}_{0}\mathbf{W}_{B}^{T}$ of process $\{{S}^{(0|{\Pi})}_{t}\}$, process $\{S^{(RP|\Pi)}_{t}\}$ is a martingale of form (B.3) plus the extra drift $R^{B}_{t}dt=-{d}(RP^{B}(t))$ that stems from $B$-sure risk-pricing:
$dS^{(RP|\Pi)}_{t}=\big{[}(\mathbf{1}_{+}-\Pi^{+}_{t})\check{r}^{Z\Delta}_{t}dt+\sigma^{Z}_{t}dw^{Z}_{t}\big{]}+S^{\Delta}(t)(\sigma^{\Pi}_{t})^{2}\big{[}(\mathbf{1}_{+}-\Pi^{+}_{t})(\sigma^{l}_{t})^{2}dt+\sigma^{l}_{t}dw^{\prime}_{t}\big{]}+R^{(\cdot)}_{t}dt,$ (B.7)
@@
where $\check{r}^{Z\Delta}_{t}dt:=-{d}{}S^{\Delta}(t)$, with $S^{\Delta}(t)\equiv Y^{\Delta}(t)-RP^{+}(t)+RP^{-}(t)$. As such, it is viable so long as this extra drift $\{R^{B}_{t}dt\}$ can be ’absorbed’ into the known $(\Pi^{B}_{0}\mathbf{W}_{B}^{T})$-martingale $\{d{S}^{(0|{\Pi})}_{t}\}$, that is, whenever
@@
is square-integrable up to horizon $T$ (Novikov’s Condition), where, for convenience, we have written $(\sigma^{l}_{t})^{2}=(\sigma^{lD}_{t})^{2}(1+\text{F}^{Z/D}_{t})$, with $\text{F}^{Z/D}_{t}:=(\sigma^{lZ}_{t})^{2}/(\sigma^{lD}_{t})^{2}$.
@@ C THE MOMENTUM AND LOW-RISK FORMULAE
For the assets of Section 3.2-3.4, all the distributions required stem from the underlying log-LR process $\{l^{1/{\text{{\bf{0}}}}}_{t}\}$, of parameter $\sigma^{l}$, for which log-LR levels $l^{1/{\text{{\bf{0}}}}}_{t}$ follow the Normal distribution:
@@
The distributions of RNE beliefs $\{\Pi^{1}_{t}\}$ and so of prices ((38)) then derive via change-of-variable:
@@
for $\Pi^{1}_{t}=v$ given $l^{1/{\text{{\bf{0}}}}}_{t}=l$, with $dl=(v\underline{v})^{-1}dv$ ((3)) and $O_{f}[\Pi^{1}_{0\pm}]=O_{f}[\mathbf{p}^{1}_{0}]/(\rho K^{\pm 1})$ ((39)) given $sign[1{\text{{\bf{0}}}}]=\pm$. The $H$-variables above are inferential milestones: $H_{\Pi\pm}$ is the log-LR hurdle for event $\{\Pi^{1}_{t}\geq\frac{1}{2}\}$, and $H_{\mathbf{p}}$, for $\{\mathbf{p}^{1}_{t}\geq\frac{1}{2}\}$.
@@ C.1 Tracking Inferential Progress and Bias Dominance
The degree of $B$-certainty as data accumulate may be assessed in the usual way given Normal distributions. The objective degree is tracked by some $|C^{\mathbf{p}}_{t}|$, where, with $t_{\mathbf{p}}:=2H_{\mathbf{p}}/(\sigma^{l})^{2}$,
@@
Certainty is high for $\sigma^{l}_{[t]}\ll 1$ (little data) or $\sigma^{l}_{[t]}\gg 1$ (lots of data), with $|C^{\mathbf{p}}_{t}|$ bottoming at 00 if and when $H_{\mathbf{p}}=\mu^{l}_{[t]}(B)$, that is, if and when $t=(-1)^{B+1}t_{\mathbf{p}}$ has a solution.
@@
[…]
Bias creates a ’burden of proof’ $t_{\rho}$, and $t_{\rho}/t$ tracks how it is overcome. We consider cases of positive objective hurdle $H_{\mathbf{p}}>0$ and bias domination $H_{\Pi\pm}-H_{\mathbf{p}}\gg 0$, so $t_{\mathbf{p}}>0$ and $t_{\rho}\gg t_{K}$, where ’objective burden’ $t_{\mathbf{p}}$ is presumed ’non-extreme’ for non-trivial inference/excess.
@@ C.2 The Window of Opportunity
Consider for example profitable events $\{O_{f}[{\Pi}^{1}_{t}]=\rho^{-1}\}$ or $\{O_{f}[{\Pi}^{1}_{t\pm u}]=\rho^{-1}\}$ ((44)), with $t$ fixed and inside the above window, and $u>0$ large enough for the second event to be outside. Their respective probability densities are $\propto(t)^{-1/2}e^{-|C^{\mathbf{p}}_{t}|^{2}/2}$ and $\propto(t\pm u)^{-1/2}e^{-|C^{\mathbf{p}}_{(t\pm u)}|^{2}/2}$. Thus, in-window events dominate, with a log-LR of $\propto\big{(}\log[1\pm u/t]+|C^{\mathbf{p}}_{(t\pm u)}|^{2}-|C^{\mathbf{p}}_{t}|^{2}\big{)}\gg 1$, which diverges as $(1-u/t)\to 0$ or $(1+u/t)\to\infty$, with $t$ fixed.
@@ C.3 Momentum and Volatility Mixtures
$\mathbb{P}_{t}(\Pi^{1}_{t}=v)\mathbb{P}_{t}(\pm|\text{ }\Pi^{1}_{t}=v)\equiv\mathbb{P}_{t}(\pm)\mathbb{P}_{t}(\Pi^{1}_{t}=v|\text{ }\pm),$ (C.9)
with $\mathbb{P}_{t}(\Pi^{1}_{t}=v|\text{ }\pm)$ from Normal distribution (C.1) via (C.2). The likelihood-ratio of $\{\Pi^{1}_{t}=\underline{v}\}$ vs $\{\Pi^{1}_{t}=v\}$, $v\in(0,\frac{1}{2}]$, given the $sign[1{\text{{\bf{0}}}}]=\pm$ of potential change, is:
@@
Note that $\{\Pi^{1}_{t}=v\}$-events dominate as long as bias does (i.e. $t_{\rho}/{t}\gg 1$) regardless; its dependence on actual $B$-outcomes is weak.
To compute (42) we also need the mix-ratio function $M_{t|v}:={\mathbb{P}_{t}(+|v)}/{\mathbb{P}_{t}(-|v)}$ of the event-set $\{\Pi^{1}_{t}=v\}$, $v\in(0,1)$, between member-events with $\{sign[1{\text{{\bf{0}}}}]=+\}$ and with $\{sign[1{\text{{\bf{0}}}}]=-\}$.
By (C.9) we have $M_{t|v}={\mathbb{P}_{t}(v|+)}/{\mathbb{P}_{t}(v|-)}$, since $\mathbb{P}_{t}(\pm)=\frac{1}{2}$ by setup. Hence:
$M_{t|v}(B)$ $=(\frac{\rho v}{\underline{v}})^{-t_{K}/t}\cdot K^{-(t_{\mathbf{p}}/t)-(-1)^{B}},$ (C.11)
@@
Note the source of conditionality: risk-pricing $K$; if $K=1$, conditional mixes equal the unconditional mix; if $K\in(1,2)$, they are perturbations of the unconditional one.
@@
Both mix-ratios (C.11-C.12) are monotone declining, and peak-volatility ($v=\frac{1}{2}$) brings $M_{t|\frac{1}{2}}=M_{t|\frac{1}{2}\frac{1}{2}}\approx\rho^{-t_{K}/t}=\exp[{-(t_{\rho}/t)\log K}]$. That is, for $\rho\gg 1$, in the window (C.8) of opportunity, when trading events are most abundant, the observed mix at high volatility can be highly negative vs the unconditional level. Mix returns to its unconditional state as $t\to\infty$ when model-risk resolves. Any uneven unconditional background merely introduces a constant factor in (C.11-C.12); all above statements remain valid.
@@ C.4 Peak-Reward Location and Size for Volatility-Conditioned Trading
Further, given (C.10) and under condition (C.8), the $\{\Pi^{1}_{t}=\underline{v}\}$-contributions to (42) are no more than $\mathcal{O}(\rho^{-t_{\rho}/{t}})$. That is, the leading order solution (44) is not affected by these contributions while $t_{\rho}/t\gg 1$, i.e. while bias dominates.
@@
For uneven unconditional-mix, $\mathbb{P}_{t}(\pm)\not=\frac{1}{2}$, the same applies. Compared to even solution (44), its peak moves left (right), and size, down (up), for unevenness that tilts negatively (positively); peak-location is capped by $v^{+mo}_{max}$ ((43)), and peak-size, by $rp^{+mo}_{max}$ ((43)).
Without uncertainty or bias, $\rho=1$, both the $\{\Pi^{1}_{t}=v\}$- and $\{\Pi^{1}_{t}=\underline{v}\}$-part of (42) are invariant under $v\leftrightarrow\underline{v}$ switching ((41) and (C.10-C.12)), so their derivatives both vanish at $v=\frac{1}{2}$, thus confirming ’high risk high reward’ under classical conditions.

### v6

@@ B PROOF OF PROPOSITION 1
This has solutions of the form $\{A^{+}_{t}\}=\{\Pi^{+}_{t}\}$ only, where $\{\Pi^{+}_{t}\}$ is any model-risk inference based on any measure $\Pi^{B}_{0}W_{B}^{Z}|_{T}$, $\Pi^{B}_{0}\sim\pi^{B}_{0}$.
@@
Adding $B$-conditionally independent data $\{D_{t}\}$ under law $W_{B}^{D}|_{T}$, the argument above applies unchanged to the dynamics of $\{dS^{(0|A)}_{t}\}$, (B.1-B.2), under the reference belief ${\pi}^{B}_{0}\mathbf{W}_{B}^{T}$ and some RNE measure $\hat{\pi}^{B}_{0}\hat{\mathbf{W}}_{B}^{T}$ such that (B.4) holds and $d\hat{\mathbf{W}}^{T}_{B,t}=\hat{r}_{B,\text{ }t}dt+d\mathbf{W}^{T}_{-,t}$, $\hat{{r}}_{B,\text{ }t}dt=-d\hat{{RP}}^{B}_{t}$. To conclude, the RNE measure $\{S^{(0|A)}_{t}\}$ must have the canonical form: $\Pi^{B}_{0}\mathbf{W}_{B}^{T}$, $\Pi^{B}_{0}\sim\pi^{B}_{0}$.
Part II. The viability of canonical price process $\{S^{(RP|{\Pi})}_{t}\}=\{{S}^{(0|{\Pi})}_{t}\}-\{\Sigma_{B}{RP}^{B}(t)\Pi^{B}_{t}\}$.
@@
Under the martingale measure $\Pi^{B}_{0}\mathbf{W}_{B}^{T}$ of process $\{{S}^{(0|{\Pi})}_{t}\}$, process $\{S^{(RP|\Pi)}_{t}\}$ is a martingale of form (B.3) plus the extra drift $R^{B}_{t}dt=-{d}RP^{B}(t)$ that stems from $B$-sure risk-pricing:
$dS^{(RP|\Pi)}_{t}=\big{[}(\mathbf{1}_{+}-\Pi^{+}_{t})\cdot\check{r}^{Z\Delta}_{t}dt+\sigma^{Z}_{t}dw^{Z}_{t}\big{]}+S^{\Delta}(t)(\sigma^{\Pi}_{t})^{2}\big{[}(\mathbf{1}_{+}-\Pi^{+}_{t})\cdot(\sigma^{l}_{t})^{2}dt+\sigma^{l}_{t}dw_{t}\big{]}+R^{(\cdot)}_{t}dt,$ (B.7)
@@
where $\check{r}^{Z\Delta}_{t}dt:=-{d}{}S^{\Delta}(t)$, with $S^{\Delta}(t)\equiv Y^{\Delta}(t)-RP^{+}(t)+RP^{-}(t)$ (by (12-14). As such, it is viable if drift $\{R^{B}_{t}dt\}$ can be ’absorbed’ into the known $(\Pi^{B}_{0}\mathbf{W}_{B}^{T})$-martingale $\{d{S}^{(0|{\Pi})}_{t}\}$, that is, whenever
@@
is square-integrable up to horizon $T$ (Novikov’s Condition), where $(\sigma^{l}_{t})^{2}\equiv(\sigma^{lD}_{t})^{2}(1+\text{F}^{Z/D}_{t})$ is written in terms of the relative variance $\text{F}^{Z/D}_{t}:=(\sigma^{lZ}_{t})^{2}/(\sigma^{lD}_{t})^{2}$ of data $\{Z_{t}\}$ to $\{D_{t}\}$.
@@ C THE MOMENTUM AND LOW-RISK FORMULAE
For the assets of Section 3.2-3.3, all the distributions required stem from the underlying log-LR process $\{l^{1/{\text{{\bf{0}}}}}_{t}\}$, of parameter $\sigma^{l}$, for which log-LR levels $l^{1/{\text{{\bf{0}}}}}_{t}$ follow the Normal distribution:
@@
The distributions of RNE beliefs $\{\Pi^{1}_{t}\}$ and so of prices (37) then derive via (3) by change of variable: for $\Pi^{1}_{t}=v$ given $l^{1/{\text{{\bf{0}}}}}_{t}=l$, we have $dl=(v\underline{v})^{-1}dv$, with:
@@
where recall (39), $O_{f}[\Pi^{1}_{0\pm}]=O_{f}[\mathbf{p}^{1}_{0}]/(\rho K^{\pm 1})$ given $sign[1{\text{{\bf{0}}}}]=\pm$. The $H$-variables above are inferential milestones: $H_{\Pi\pm}$ is the log-LR hurdle for event $\{\Pi^{1}_{t}\geq\frac{1}{2}\}$, and $H_{\mathbf{p}}$, for $\{\mathbf{p}^{1}_{t}\geq\frac{1}{2}\}$.
@@ C.1 Tracking Inferential Progress and Bias Dominance
The degree of $B$-certainty as data accumulate may be assessed in the standard way for Normal distributions. Its objective level is tracked by some $|C^{\mathbf{p}}_{t}|$, where, with $t_{\mathbf{p}}:=2H_{\mathbf{p}}/(\sigma^{l})^{2}$,
@@
Certainty level $|C^{\mathbf{p}}_{t}|$ is high if $\sigma^{l}_{[t]}\ll 1$ (little data) or $\sigma^{l}_{[t]}\gg 1$ (lots of data); it bottoms out at 0 if and when $H_{\mathbf{p}}=\mu^{l}_{[t]}(B)$ has a solution, that is, if and when $t=(-1)^{B+1}t_{\mathbf{p}}$ and $B=1$.
@@
where $t_{\Pi\pm}=t_{\rho}\pm t_{K}+t_{\mathbf{p}}$, with $t_{\rho}:={2\log(\rho)}/{(\sigma^{l})^{2}}>0$ and $t_{K}:={2\log K}/{(\sigma^{l})^{2}}>0$ (recall $K-1\not\gg 1$ in general and $K\in(1,2)$ in competitive markets). The pattern of $|C^{\Pi}_{t\pm}|$ parallels that of $|C^{\mathbf{p}}_{t}|$; it has 0 as a minimum if and when $t=(-1)^{B+1}t_{\Pi\pm}>0$.
Bias creates a ’burden of proof’ $t_{\rho}$, and $t_{\rho}/t$ tracks how it is overcome. We consider cases of positive objective hurdle $H_{\mathbf{p}}>0$ and bias domination $H_{\Pi\pm}-H_{\mathbf{p}}\gg 0$, so $t_{\mathbf{p}}>0$ and $t_{\rho}\gg t_{K}$, where ’objective burden’ $t_{\mathbf{p}}$ is ’non-extreme’, to exclude easy inference or profits.
@@ C.2 The Window of Opportunity
Consider profitable events $\{O_{f}[{\Pi}^{1}_{t}]=\rho^{-1}\}$ and $\{O_{f}[{\Pi}^{1}_{t\pm u}]=\rho^{-1}\}$ (see (43)), with $t$ fixed inside the window above and $u>0$ large enough for the second event to be outside the window. Their respective probability densities are $\propto(t)^{-1/2}e^{-|C^{\mathbf{p}}_{t}|^{2}/2}$ and $\propto(t\pm u)^{-1/2}e^{-|C^{\mathbf{p}}_{(t\pm u)}|^{2}/2}$. Thus, in-window events dominate, with a log-LR of $\log[1\pm u/t]+|C^{\mathbf{p}}_{(t\pm u)}|^{2}-|C^{\mathbf{p}}_{t}|^{2}\gg 1$, which diverges as $(1-u/t)\to 0$ or $(1+u/t)\to\infty$, $t$ fixed.
@@ C.3 Momentum and Volatility Mixtures
[…]
with $\mathbb{P}_{t}(\Pi^{1}_{t}=v|\text{ }\pm)$ from Normal distribution (C.1) via (C.2). The likelihood-ratio of events $\{\Pi^{1}_{t}=\underline{v}\}$ vs $\{\Pi^{1}_{t}=v\}$, $v\in(0,\frac{1}{2}]$, given the $sign[1{\text{{\bf{0}}}}]=\pm$ of potential change, is:
@@
Note that $\{\Pi^{1}_{t}=v\}$-events dominate when bias does (i.e. $t_{\rho}/{t}\gg 1$), regardless of $B$-outcomes.
To compute (41) we also need the mix-ratio function $M_{t|v}:={\mathbb{P}_{t}(+|v)}/{\mathbb{P}_{t}(-|v)}$ of the event-set $\{\Pi^{1}_{t}=v\}$, $v\in(0,1)$, between member-events with $\{sign[1{\text{{\bf{0}}}}]=+\}$ and with $\{sign[1{\text{{\bf{0}}}}]=-\}$. Given (C.9), we have $M_{t|v}={\mathbb{P}_{t}(v|+)}/{\mathbb{P}_{t}(v|-)}$, since $\mathbb{P}_{t}(\pm)=\frac{1}{2}$ by setup. Hence:
$M_{t|v}(B)$ $=(\frac{\rho v}{\underline{v}})^{-t_{K}/t}K^{-(t_{\mathbf{p}}/t)-(-1)^{B}},$ (C.11)
@@
Note the source of the conditionalities above, model-risk pricing $K$. For $K=1$, unpriced model-risk, conditional mixes equal the unconditional. For $K\in(1,2)$, competitively priced model-risk, conditional mixes are perturbations of the unconditional.
@@
Both mix-ratios (C.11-C.12) are monotone declining in $v$, and peak-volatility ($v=\frac{1}{2}$) brings $M_{t|\frac{1}{2}}=M_{t|\frac{1}{2}\frac{1}{2}}\approx\rho^{-t_{K}/t}=\exp[{-(t_{\rho}/t)\log K}]$. That is, for $\rho\gg 1$, in the window (C.8) of opportunity, when trading events are most abundant, the observed mix at high volatility can be highly negative vs the unconditional level. Mix reverts to its unconditional level as $t\to\infty$ when model-risk resolves. Any uneven unconditional background merely introduces a constant factor in (C.11-C.12); all above statements remain valid.
@@ C.4 Peak-Reward Location and Size for Volatility-Conditioned Trading
Further, given (C.10) and under condition (C.8), the $\{\Pi^{1}_{t}=\underline{v}\}$-contributions to (41) are no more than $\mathcal{O}(\rho^{-t_{\rho}/{t}})$. That is, the leading order solution (43) is not affected by these contributions whenever $t_{\rho}/t\gg 1$, i.e. while bias dominates.
@@
For uneven unconditional-mix, $\mathbb{P}_{t}(\pm)\not=\frac{1}{2}$, the same features apply. For unevenness that tilts negatively (positively), the location of peak-excess moves left (right) vs the even solution (43), and its size, down (up).
Without uncertainty or bias, $\rho=1$, both the $\{\Pi^{1}_{t}=v\}$- and $\{\Pi^{1}_{t}=\underline{v}\}$-contributions to (41) are invariant under $v\leftrightarrow\underline{v}$ switching (see (40) and (C.10-C.12)), so their derivatives both vanish at $v=\frac{1}{2}$, thus confirming peak-volatility as the location of peak-reward and ’high risk high reward’ under classical conditions.